OpenStudy (anonymous):

Find all real solutions x3 +x+1=x+9

OpenStudy (anonymous):

x^3 +x -x +1 -9 =0 x^3 -8 = 0 x^3 = 8 x = 2

OpenStudy (anonymous):

yep